OpenStudy (anonymous):

Find the x-coordinates of all points on the curve f(x) = sin 2x − 2 sin x at which the tangent line is horizontal. (Enter your answers as a comma-separated list. Use n to represent any integer.)

OpenStudy (anonymous):

is it \(f(x)=\sin(2x)-2\sin(x)\) ?

OpenStudy (anonymous):

I suppose so. The question just has sin 2x - 2 sin x, but that is how I'd read it yeah

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

\[f'(x)=2\cos(2x)-2\cos(x)\]

OpenStudy (anonymous):

\[2\cos(2x)-2\cos(x)=0\] \[\cos(2x)=\cos(x)\]

OpenStudy (anonymous):

hmm how to solve any even multiple of \(\pi\) works, but i am sure there are more maybe start with \[\cos(2x)-\cos(x)=0\] then rewrite as \[2\cos^2(x)-\cos(x)-1=0\] and solve the quadratic for cosine

OpenStudy (anonymous):

Pretty lost. How do I know if the tangent is horizontal or even where it is?

OpenStudy (anonymous):

tangent horizontal = derivative is zero

OpenStudy (anonymous):

solve \[2\cos^2(x)-\cos(x)-1=0\] \[(2\cos(x)+1)(\cos(x)-1)=0\]\[\cos(x)=-\frac{1}{2}\] or \[\cos(x)=1\]

OpenStudy (anonymous):

I don't get it. How many values should the answer be? is the answer where those equal 0?